Determine the complex number z which satisfies the equations |z + 3i| = |z + 5 − 2i|
and |z − 4i| = |z + 2i| simultaneously.
1
Expert's answer
2020-05-19T08:49:28-0400

We have given that a complex number "z" satisfy these two equation

"|z − 4i| = |z + 2i| \hspace{1cm}(1)\\
|z + 3i| = |z + 5 − 2i|\hspace{1cm}(2)"

Let's assume the solution is "z=z_0=x_0+y_0" ,thus equation (1) will be,


"|z_0-4i|=|z_0+2i|\\
\implies |z_0-4i|^2=|z_0+2i|^2\\
\implies (z_0-4i)(\overline{z_0}+4i)=(z_0+2i)(\overline{z_0}-2i) \hspace{1cm}(\because |z|^2=z\overline{z})\\
\implies 6i(z_0-\overline{z_0})=-12 \hspace{1cm}(\because z_0=x_0+y_0)\\
\implies -12y_0=-12 \implies y_0=1"

Hence,

"z_0=x_0+i \hspace{1cm}(3)"

And similarly, from equation (2) we get,


"|z_0 + 3i| = |z_0+ 5 − 2i|\\
\implies |(x_0+i) + 3i|^2 = |(x_0+i) + 5 − 2i|^2 \hspace{1cm}(\because z_0=x_0+i)\\
\implies |x_0 + 4i|^2 = |(x_0 + 5) − i|^2\\
\implies x_{0}^2 + 16 = (x_0 + 5)^2 + 1\\
\implies 10x_0=-10 \implies x_0=-1"

Thus,

"z_0=-1+i"

Therefore, required complex number

"z=z_0=-1+i"

which satisfy the equation "(1) \: \& (2)". Hence we are done.